please help me solve
3a+b=1........(i)
2a+3b=8.......(ii)

Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

please help me solve
3a+b=1........(i)
2a+3b=8.......(ii)
------------
Multiply the 1st eqn by 3, then subtract.

Answer by addingup(3677)

3a+b = 1
2a+3b = 8
I'm not sure what you want to do here, you need to specify what it is that you need.
I'll solve by elimination and hopefully it helps you.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
3a+b = 1 multiply both sides times -3 and add to the next equation:
-9a+(-3b) = -3
+
2a+3b = 8
---------------------
-7a 0b = 5
-7a = 5
a = -5/7
Now find b by substituting a with -5/7
3(-5/7)+b = 1
-15/7 + b = 1
Rewrite like this:
b + -15/7 = 1
b = 1 + 15/7
b = 22/7
~~~~~~~~~~~~~~~~~~~~
Check:
3(-5/7) + (22/7) = 1
-15/7 + 22/7 = 7/7 = 1 Well, this one’s correct
2(-5/7) + 3(22/7) = 8
-10/7 + 66/7 = 56/7 = 8 How do you know, this one’s also correct ;-)